334. A manufacturer claims that 10% of women using the "pill" suffer from side effects. The Federal Drug Administration (FDA) believes that the manufacturer's claim is too low and decides to test the manufacturer's claim at α = 5 %. A random sample of 900 women who use the manufacturer's pill shows that 103 suffer side effects. The FDA's null hypothesis is:
Answer #1

Solution :

Given that,

Po = 0.10

1 - Po = 0.90

n = 900

x = 103

Level of significance = \alpha = 0.05

Point estimate = sample proportion = \hat p = x / n = 0.114

This a left (One) tailed test.

The null and alternative hypothesis is,

Ho: p = 0.10

Ha: p < 0.10

Test statistics

z = (\hat p - Po ) / \sqrt{} Po*(1-Po) / n

= ( 0.114- 0.10) / \sqrt{} (0.10*0.90) / 900

= 1.444

P-value = P(Z < z )

= P(Z < 1.444 )

= 0.9257

The p-value is p = 0.9257, and since p = 0.9257 > 0.05, it is concluded that the null hypothesis is fail to rejected.